### v1

# Quantum mechanics, non-locality, and the space discreteness hypothesis

Abstract.
[…]
1. Introduction
Quantum mechanics (QM) is highly successful in describing the behavior of matter and energy at the atomic and subatomic levels, which has driven the development of technologies that have reshaped human society. However, there is a consensus that QM has significant flaws; if the theory is not wrong, it is at least incomplete. One major issue is the measurement problem, specifically the collapse of the wave function during measurement. Other issues are the incompatibility with general relativity and the existence of several interpretations of quantum mechanics.
Non-locality of QM means that the components of a quantum system may continue influence each other, even where they are well-separated and out of speed-of-light contact. It was first pointed out by Einstein, Podolsky and Rosen (EPR) in 1935, [1]. EPR pointed out that non-locality implies that the standard formalism of QM is incomplete or wrong. In the beginning of the 1970s, Freedman and Clauser conducted experiments to test  the Bell-inequality violation, and other aspects of entangled quantum systems. These experiments showed that both QM and the underlying reality it describes are intrinsically non-local. Einstein’s spooky actions-at-a-distance really happened in the physical world. For an in-depth discussion, the reader may consult [2]-[5], and the references therein.
In the theory of relativity (TR), the use of manifolds that locally look like $\mathbb{R}\times\mathbb{R}^{3}$ to describe the universe at large scales has been highly successful. We argue that the clash between QM and the TR emerges when the model of space-time $\mathbb{R}\times\mathbb{R}^{3}$ is used to understand quantum phenomena. At the simplest possible level, causality emerges from the total order of $\mathbb{R}$, and locality from the fact that the standard derivatives define local operators, which in turn do not allow spooky action at a distance.
The problem described above is know since 1920’s; the creators of the QM were convinced that the classical space-time continuum must abandoned on the elementary quantum scale. Some them believed that the traditional concept of space-time of macroscopic physics cannot simply be transferred to quantum physics. But all the efforts to construct physical theories with discrete space-times failed, [6].
In the 1930s, Bronstein showed that general relativity and quantum mechanics imply that the uncertainty $\Delta x$ of any length measurement is greater than or equal to the Planck length, [7]. The interpretation of this inequality is that there are no intervals, just points below the Planck length. We argue that this last assertion can be reformulated in precise mathematical form as: at the Planck length, the physical space can be modeled as a completely disconnected space $\mathcal{X}$. In this framework, there is no continuous world line joining two different points in space $\mathcal{X}$; consequently, the assumption that $\mathcal{X}$ is a model of space is entirely incompatible with the theory of relativity. In this paper,  $\mathcal{X}$ is the mathematical realization of the hypothesis of the discreteness of space at small distances. The naive idea of using a lattice in $\mathbb{R}^{3}$ as a model for the space discreteness is not useful at all.
We propose $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{3}$ as a model of the space-time (or more generally, $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{N}$ as a model for the configuration space). The advantage of this model is that it contains regions homeomorphic to $\mathbb{R}\times\mathbb{R}^{3}$ and to $\mathbb{R}\times\mathcal{X}^{3}$. In the first type of regions, causal and local theories like the relativistic QM and TR can be formulated. In contrast, in the second type of regions, acausal and non-local theories can be formulated. In this framework,  QM (in sense of Dirac-von Neumann) on the Hilbert spaces $L^{2}\left(\left(\mathbb{R}\times\mathcal{X}\right)^{N}\right)$ can be rigorously formulated, [8]-[11]. A rigorous mathematical study of QM on Hilbert spaces of type $L^{2}\left(\left(\mathbb{R}\times\mathcal{X}\right)^{N}\right)$ is possible assuming suitable technical conditions, but this choice is quite involved. Here, we take $N=1$, $\mathbb{Q}_{p}$, the field of $p$-adic numbers; $\mathbb{Q}_{p}$ is a paramount example of a completely disconnected space, with a very rich mathematical structure. In the following paragraphs, we provide an argument showing the physical relevance of the chosen model.
In the Dirac-von Neumann formulation of QM, the states of a closed quantum system are vectors of an abstract complex Hilbert space $\mathcal{H}$, and the observables correspond to linear self-adjoint operators in $\mathcal{H}$. In $p$-adic QM, $\mathcal{H}=L^{2}(\mathbb{Q}_{p}^{N})$ or $\mathcal{H}=L^{2}(U)$, where $U\subset\mathbb{Q}_{p}^{N}$. This choice implies that the position vectors are $\mathbb{Q}_{p}^{N}$, while the time is a real variable. Using $\mathbb{R}\times\mathbb{Q}_{p}^{3}$ as a model of space-time, implies assuming that the space ($\mathbb{Q}_{p}^{3}$) has a discrete nature: there are no continuous word lines joining two different points in the space. We warn the reader that there are several different types of $p$-adic QM; for instance, if the time is assumed to be a $p$-adic variable, the QM obtained radically differs from the one considered here. To the best of our knowledge, $p$-adic QM (with $p$-adic time) started in the 1980s under the influence of Vladimirov and Volovich, [14]. The literature about $p$-adic QM is pervasive, and here we cited just a few works, [12]-[25].
A fundamental problem of $p$-adic quantum mechanics was the lack of physical interpretation. In [21], see also [24]-[25], the author showed that a large class of $p$-adic Schrödinger equations is the scaling limit of certain continuous-time quantum Markov chains (CTQMCs). This construction includes, as a particular case, the continuous time quantum walks (CTQWs) constructed using adjacency matrices, [26]-[29]. The CTQWs are a fundamental tool in the development of quantum algorithms. The connection between $p$-adic QM and CTQWs shows that $p$-adic QM has a physical meaning. $p$-Adic QM is a non-local theory (the Hamiltonians used are non-local operators), and consequently, spooky actions at a distance are allowed. The paradigm asserting that the universe in not locally real, implies that $p$-adic QM allows realism.
The non-locality tests are formulated using QM on $\mathcal{H}=\mathbb{C}^{N}$, see, e.g., [3], [30]- [33], and the references therein. In Section 3, we construct an $N$-dimensional Hilbert space $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$, spanned by a set of functions whose supports are contained in the unit ball $\mathbb{Z}_{p}$. Since $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ is isomorphic to $\mathbb{C}^{N}$ as Hilbert spaces, the non-locality tests can be recast as assertions in $p$-adic QM on the Hilbert space $L^{2}(\mathcal{\chi}_{N}(\mathbb{Z}_{p}))\subset L^{2}(\mathbb{Z}_{p})$. In addition, any Hamiltonian on $\mathbb{C}^{N}\simeq\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ (a Hermitian matrix) has an extension as a non-local operator on $L^{2}(\mathbb{Z}_{p})$. In our view, this argument, and the connection with quantum computing, establish the relevance of the choice $\mathcal{X}=\mathbb{Q}_{p}$.
Our first result is a theoretical framework, where QM is a non-local theory admitting realism. An important step is to pass from a phenomenological description of non-locality, as the one given in the second paragraph above, to a mathematical one. We propose understanding non-locality in QM as the emergence of the use of non-local operators (Hamiltonians) in quantum models. These types of operators naturally allow spooky action at a distance, and then this approach is consistent with the phenomenological one. However, the use of local Hamiltonians (which do not enable spooky action at a distance) contradicts the experimental confirmation of non-locality.
[…]
$i\frac{\partial}{\partial t}\Psi_{\infty}\left(x_{\infty},t\right)=\boldsymbol{H}_{\infty}\Psi_{\infty}\left(x_{\infty},t\right),$
[…]
$i\frac{\partial}{\partial t}\Psi_{p}\left(x_{p},t\right)=\boldsymbol{H}_{p}\Psi_{p}\left(x_{p},t\right),$
where $\boldsymbol{H}_{p}:Dom(\boldsymbol{H}_{p})\rightarrow L^{2}\left(\mathbb{Q}_{p}\right)$ is a linear, non-local Hamiltonian densely defined in $L^{2}\left(\mathbb{Q}_{p}\right)$. Now, by using the ansatz
$\Psi\left(x_{\infty},x_{p},t\right)=\Psi_{\infty}\left(x_{\infty},t\right)\Psi_{p}\left(x_{p},t\right),$
[…]
(1.1) $i\frac{\partial}{\partial t}\Psi\left(x_{\infty},x_{p},t\right)=\left(\boldsymbol{H}_{p}+\boldsymbol{H}_{\infty}\right)\Psi\left(x_{\infty},x_{p},t\right),$
[…]
In Section 6, we apply the theoretical framework developed to the measurement problem. In this framework, the collapse of the wavefunction is a consequence of the geometry of the space-time. The collapse mechanism resembles the one proposed by Ghirardi, Ramini, and Weber (GRW); see, e.g., [33]-[31], but, there are three major differences. First, the wavefunction is not a physical entity; second, we do not need to introduce new physical constants. Third, the Schrödinger equation (1.1) described the collapse of the wavefunction, so, it is not necessary to introduce a new dynamical equation.
[…]
The model introduced here opens many questions and controversies. In Section 8, we discuss briefly some matters like Einstein’s causality, the no-communication theorem, and the nature of space-time at small distances, etc.
2. The space discreteness hypothesis
2.1. The Bronstein inequality
Since the 1920s, Born, Einstein, Jordan, Pauli, among others, were convinced that the classical space-time continuum must abandoned on the elementary quantum scale. The reader may consult [6] for an in-depth historical review. For instance, according to [6], Einstein and Born believed that the traditional concept of space-time of macroscopic physics cannot simply be transferred to quantum physics. But all the efforts (especially from Born and Jordan) to construct physical theories with discrete space-times failed.
In the 1930s, Bronstein showed that general relativity and quantum mechanics imply that the uncertainty $\Delta x$ of any length measurement satisfies
(2.1) $\Delta x\geq L_{\text{Planck}}:=\sqrt{\frac{\hbar G}{c^{3}}},$
where $L_{\text{Planck}}$ is the Planck length ($L_{\text{Planck}}\approx 10^{-33}$ cmcm), [7]. The standard interpretation of Bronstein’s inequality is that below the Planck length, there are no intervals, just points. The intervals are the basic examples of connected sets; such a set is the mathematical incarnation of an infinitely divisible object. Then, below the Planck length, the topology of space only allows trivial (points) connected subsets. This can be recast as below the Planck length, the space is a totally (or completely) disconnected topological space.
We denote by $\mathcal{X}$ a totally disconnected space. Notice $\mathbb{R}^{N}$ is connected space. Consider a word line (a continuous curve) in $\mathcal{X}$:
$\begin{array}[c]{cccc}x:&\left[a,b\right]&\rightarrow&\mathcal{X}\\ &t&\rightarrow&x\left(t\right).\end{array}$
We recall that continuous mappings preserve connectedness. Then, the image $x\left(\left[a,b\right]\right)$ must be a connected subset of $\mathcal{X}$, since the only connected subsets are points; the curve $x\left(t\right)$ is constant. In other words, given two different points in $\mathcal{X}$, there is no word line joining them. The motion in $\mathcal{X}$ is a sequence of jumps. Consequently, we replace the space-time model $\mathbb{R}\times\mathbb{R}^{3}$ by $\mathbb{R}\times\mathcal{X}^{3}$, i.e., if we assume the discreteness of space, then we lose the compatibility with special and general relativity.
2.2. Macroscopic versus microscopic
We divide the space-time into two realms: the macroscopic and the microscopic. In the first one, the model of the space-time is $\mathbb{R}\times\mathbb{R}^{3}$, while in the second one, it is $\mathbb{R}\times\mathcal{X}^{3}$, where $\mathcal{X}$ is a completely disconnected topological space. The connectedness of a topological space is a topological property, which means it remains invariant under continuous transformations (homeomorphisms). Therefore, $\mathcal{X}$ is not topologically equivalent to $\mathbb{R}$, nor a proper subset of $\mathbb{R}$.
The formulation of models using $\mathbb{R}\times\mathcal{X}^{N}$ as a configuration space requires that algebraic operations, like sum and product, be defined on $\mathcal{X}$. A natural candidate for $\mathcal{X}$ is a $K$-analytic Lie group, where $K$ is a non-Archimedean field, [39]. It is not possible to assume a sharp boundary between the macroscopic and microscopic realms, and, in addition, there is no topologically faithful copy of $\mathcal{X}$ in $\mathbb{R}$; for this reason, we propose
$\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{3}$
as a model of space-time. Due to mathematical convenience, we use the word space-time in an extended way to mean configuration space. Then, we propose using configuration spaces of type $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{N}$.
2.3. The $p$-adic space
In this section, we fix the notation and collect some basic results on $p$-adic analysis that we use in this paper. For a detailed exposition on $p$-adic analysis, the reader may consult [15], [40], [41], [42]. For a discussion on applications of $p$-adic analysis to physics, biology, computing, and other areas, the reader may consult [19], [22], [25], [43], [44], [45], [46], [48], and the references therein.
From now on, we use $p$ to denote a fixed prime number. Any non-zero $p$-adic number $x$ has a unique expansion of the form
(2.2) $x=x_{-k}p^{-k}+x_{-k+1}p^{-k+1}+\ldots+x_{0}+x_{1}p+\ldots,\text{ }$
with $x_{-k}\neq 0$, where $k$ is an integer, and the $x_{j}$s are numbers from the set $\left\{0,1,\ldots,p-1\right\}$. The set of all possible sequences of the form (2.2) constitutes the field of $p$-adic numbers $\mathbb{Q}_{p}$. There are natural field operations, sum and multiplication, on series of form (2.2). There is also a norm in $\mathbb{Q}_{p}$ defined as $\left|x\right|_{p}=p^{-ord(x)}$, where $ord_{p}(x)=ord(x)=k$, for a nonzero $p$-adic number $x$. By definition $ord(0)=\infty$. The field of $p$-adic numbers with the distance induced by $\left|\cdot\right|_{p}$ is a complete ultrametric space. The ultrametric property refers to the fact that $\left|x-y\right|_{p}\leq\max\left\{\left|x-z\right|_{p},\left|z-y\right|_{p}\right\}$ for any $x$, $y$, $z$ in $\mathbb{Q}_{p}$. The $p$-adic integers which are sequences of form (2.2) with $k\geq 0$. All these sequences constitute the unit ball $\mathbb{Z}_{p}$. The unit ball is an infinite rooted tree with fractal structure. As a topological space $\mathbb{Q}_{p}$ is homeomorphic to a Cantor-like subset of the real line, see, e.g., [15], [42]. $\mathbb{Q}_{p}$ is a paramount example of a completely disconnected space, with a very rich mathematical structure.
A function $\varphi:\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is called locally constant, if for any $a\in\mathbb{Q}_{p}$, there is an integer $l=l(a)$, such that
$\varphi\left(a+x\right)=\varphi\left(a\right)\text{ for any }|x|_{p}\leq p^{l}.$
The set of functions for which $l=l\left(\varphi\right)$ depends only on $\varphi$ form a $\mathbb{C}$-vector space denoted as $\mathcal{U}_{loc}\left(\mathbb{Q}_{p}\right)$. We call $l\left(\varphi\right)$ the exponent of local constancy. If $\varphi\in\mathcal{U}_{loc}\left(\mathbb{Q}_{p}\right)$ has compact support, we say that $\varphi$ is a test function. We denote by $\mathcal{D}(\mathbb{Q}_{p})$ the complex vector space of test functions. There is a natural integration theory so that $\int_{\mathbb{Q}_{p}}\varphi\left(x\right)dx$ gives a well-defined complex number. The measure $dx$ is the Haar measure of $\mathbb{Q}_{p}$.
Let $U\subset\mathbb{Q}_{p}$ be an open subset. We denote by $\mathcal{D}(U)$ denotes the $\mathbb{C}$-vector space of test functions with supports contained in $U$; then, $\mathcal{D}(U)$ is dense in
$L^{\rho}(U)=\left\{\varphi:(U)\rightarrow\mathbb{C};\left\|\varphi\right\|_{\rho}=\left\{{\displaystyle\int\limits_{U}}\left|\varphi\left(x\right)\right|^{\rho}dx\right\}^{\frac{1}{\rho}}<\infty\right\},$
for $1\leq\rho<\infty$, see, e.g., [42, Section 4.3].
We now take $U=\mathbb{Z}_{p}$ and $\rho=2$. The density $\mathcal{D}(\mathbb{Z}_{p})$ in $L^{2}\left(\mathbb{Z}_{p}\right)$ means that given $f\in L^{2}\left(\mathbb{Z}_{p}\right)$ and $\epsilon>0$, there exist a nonnegative integer $l$, and
$\varphi^{\left(l\right)}\left(x\right)=\sum_{I\in G_{l}}\varphi_{I,l}\Omega\left(p^{l}\left|x-I\right|_{p}\right)\in\mathcal{D}(\mathbb{Z}_{p}),$
such that $\left\|f-\varphi^{\left(l\right)}\right\|_{2}<\epsilon$. Here, $I=I_{0}+\ldots+I_{l-1}p^{l-1}$, with $I_{k}\in\left\{0,\ldots,p-1\right\}$, $\varphi_{I,l}\in\mathbb{C}$, and $\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ is the characteristic function of the ball $I+p^{l}\mathbb{Z}_{p}$. The function $\varphi^{\left(l\right)}$ is a discretization of $f$; the functions $\varphi^{\left(l\right)}$ constitute a $\mathbb{C}$-vector space $\mathcal{D}_{l}(\mathbb{Z}_{p})$, which is isometric to the finite-dimensional Hilbert space $\mathbb{C}^{p^{l}}$. The functions $\left\{p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-I\right|_{p}\right)\right\}_{I\in G_{l}}$ form an orthonormal basis, indeed
$\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-I\right|_{p}\right),p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-J\right|_{p}\right)\right\rangle=p^{l}{\displaystyle\int\limits_{\mathbb{Z}_{p}}}\Omega\left(p^{l}\left|x-I\right|_{p}\right)\overline{\Omega\left(p^{l}\left|x-J\right|_{p}\right)}dx$
$=p^{l}{\displaystyle\int\limits_{\left(I+p^{l}\mathbb{Z}_{p}\right)\cap\left(J+p^{l}\mathbb{Z}_{p}\right)}}dx=p^{l}\delta_{I,J}{\displaystyle\int\limits_{I+p^{l}\mathbb{Z}_{p}}}dx=\delta_{I,J},$
where $\delta_{I,J}$ denotes the Kronecker delta.
2.4. $p$-Adic QM
In the Dirac-Von Neumann formulation of QM, [10]-[11], to every isolated quantum system there is associated a separable complex Hilbert space $\mathcal{H}$ called the space of states. The states of a quantum system are described by non-zero vectors from $\mathcal{H}$. Each observable corresponds to a unique linear self-adjoint operator in $\mathcal{H}$. The most important observable of a quantum system is its energy. We denote the corresponding operator by $\boldsymbol{H}$. Let $\Psi_{0}\in\mathcal{H}$ be the state at time $t=0$ of a certain quantum system. Then at time $t$ the system is represented by the vector $\Psi\left(t\right)=\boldsymbol{U}_{t}\Psi_{0}$, where $\boldsymbol{U}_{t}=e^{-it\boldsymbol{H}}$, $t\geq 0$, is a unitary operator called the evolution operator. The vector function $\Psi\left(t\right)$ is differentiable if $\Psi\left(t\right)$ is contained in the domain $Dom(\boldsymbol{H})$ of $\boldsymbol{H}$, which happens if at $t=0$, $\Psi_{0}\in Dom(\boldsymbol{H})$, and in this case the time evolution of $\Psi\left(t\right)$ is controlled by the Schrödinger equation $i\frac{\partial}{\partial t}\Psi\left(t\right)=\boldsymbol{H}\Psi\left(t\right)$, where $i=\sqrt{-1}$ and the Planck constant is assumed to be one.
In $p$-adic QM, $\mathcal{H=}L^{2}\left(U\right)$, where $U$ is an open compact subset of $\mathbb{Q}_{p}^{N}$. Notice that the spatial variables are vectors in $\mathbb{Q}_{p}^{N}$, but the time is real variable. This last  condition is completely necessary to use the Dirac-von Neumann formalism. Now, the Hilbert spaces $L^{2}\left(\mathbb{Q}_{p}^{3}\right)$ and $L^{2}\left(\mathbb{R}^{3}\right)$ are isometric, because both have countable orthonormal bases; but the QM on $\mathcal{H=}L^{2}\left(\mathbb{R}^{3}\right)$ is a relativistic theory, but QM on $\mathcal{H=}L^{2}\left(\mathbb{Q}_{p}^{3}\right)$ is not. Then, the geometry of the configuration space changes radically the properties of the quantum models considered.
A natural ‘philosophical’ question is the existence of a universal model $\mathcal{X}$ for the space at the microscopic level. Here, we argue that in QM, the selection of a Hilbert space also includes the selection of a configuration space $\mathcal{X}$, and that the selection is made considering the particular phenomena to be modeled. This is a well established practice in QM, as Berezin and Shubin explain in their book [10]: “The question of how, for a concrete physical system, to describe the state space $\mathcal{H}$, and establish a correspondence between the observables and self-adjoint operators in $\mathcal{H}$ goes beyond a purely mathematical theory and belongs to the domain of physical practice and intuition. In every particular case this question should be handled by an expert physicist.”
2.5. Non-locality in $p$-adic QM
As discussed in the introduction, here, it is more convenient to understand non-locality as the emergence of non-local operators in physical models (the Hamiltonians are non-local). Non-local Hamiltonians allow spooky action at a distance. In the phenomenological approach, non-locality means the existence of correlations between separated particles that are stronger than what classical physics allows. This, in turn, implies that measuring a property of one particle can instantaneously influence the properties of another, even if they are far apart. This interpretation of non-locality is entirely consistent with the hypothesis of non-local Hamiltonians; however, it comes into contradiction with the hypothesis that the Hamiltonians are local operators. Here, it is relevant to recall that the locality of the Lagrangian densities is a required hypothesis in the formulation of functional integrals in field theory, [49, Section 5.1].
Along this paper, we use discrete space as a synonymous of completely disconnected space. The non-locality is natural in discrete spaces like $\mathbb{Q}_{p}$. On a completely disconnected topological space, like $\mathbb{Q}_{p}$, the locally constant functions play a central role. A function $\varphi:\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is locally constant, if $\varphi\left(x+x^{\prime}\right)-\varphi\left(x\right)=0$ for $x^{\prime}$ sufficiently small. Then, the standard notion of derivative defined as
$\lim_{x^{\prime}\rightarrow 0}\frac{\varphi\left(x+x^{\prime}\right)-\varphi\left(x\right)}{\left|x^{\prime}\right|_{p}}=0$
is not useful. However, in the $p$-adic framework, the notion of  fractional derivative is available. The simplest Hamiltonian is $\boldsymbol{D}^{\alpha}$, $\alpha>0$, the Taibleson-Vladimirov fractional, which is defined as
(2.3) $\boldsymbol{D}^{\alpha}\varphi\left(x\right)=\frac{1-p^{\alpha}}{1-p^{-\alpha-1}}{\displaystyle\int\limits_{\mathbb{Q}_{p}}}\frac{\varphi(z)-\varphi(x)}{|z-x|_{p}^{\alpha+1}}\,dz,$
for $\varphi$ a locally constant function with compact support; see, e.g., [40, Chapter 2]. To see the non-local nature of this operator, we take $\varphi\left(x\right)=1$ if $|x|_{p}\leq 1$, otherwise $\varphi\left(x\right)=0$, then
$\boldsymbol{D}^{\alpha}\varphi\left(x\right)$ $=\frac{1-p^{\alpha}}{1-p^{-\alpha-1}}\left\{{\displaystyle\int\limits_{|z|_{p}\leq 1}}\frac{\varphi(z)-\varphi(x)}{|z-x|_{p}^{\alpha+1}}\,dz+{\displaystyle\int\limits_{|z|_{p}>1}}\frac{\varphi(z)-\varphi(x)}{|z-x|_{p}^{\alpha+1}}\,dz\right\}$
$=\left\{\begin{array}[c]{lll}-\frac{1-p^{\alpha}}{1-p^{-\alpha-1}}\left(\text{ }{\displaystyle\int\limits_{|z|_{p}>1}}\frac{dz}{|z|_{p}^{\alpha+1}}\,\right)&\text{if}&|x|_{p}\leq 1\\ &&\\ \frac{1-p^{\alpha}}{1-p^{-\alpha-1}}\frac{1}{|x|_{p}^{\alpha+1}}&\text{if}&|x|_{p}>1.\end{array}\right$.
Then, the Hamiltonian $\boldsymbol{D}^{\alpha}$ allows spooky action at a distance.
2.6. Ultrametricity
Ultrametricity in physics means the emergence of ultrametric spaces in physical models. A metric space $(M,d)$ is ultrametric if  the distance $d$ satisfies $d(A,B)\leq\max\left\{d(A,C),d(C,B)\right\}$ for any three points $A$, $B,$ $C$ in $M$. Ultrametricity was discovered in the 80s by Parisi and others in the theory of spin glasses and by Frauenfelder and others in physics of proteins; see, e.g., [50], [43], see also [44, Chapter 4], and  the references therein. The $p$-adic heat equations are continuous versions (scaling limits) of master equations of complex hierarchical systems, [46], [51]-[53]. By performing a Wick rotation, in a $p$-adic heat equation, one gets a $p$-adic Schrödinger equation. This type of equations can be interpreted as  scaling limits of continuous-time quantum Markov chains, [21], [25], and the references therein.
3. $p$-Adic QM is the quantum model for quantum computing
In [21], see also [25], we showed that a large class of $p$-adic Schrödinger equations is the scaling limit of certain continuous-time quantum Markov chains (CTQMCs). Practically, a discretization of such an equation gives a CTQMC. As a practical result, we construct new types of continuous-time quantum walks (CTQWs) on graphs using two symmetric matrices. This construction includes, as a particular case, the CTQWs constructed using adjacency matrices, [26]-[27], [29]. The connection between $p$-adic QM and CTQWs show that $p$-adic QM has a physical meaning. $p$-Adic QM is a non-local theory because the Hamiltonians used are non-local operators, and consequently, spooky actions at a distance are allowed. But, this theory is not a mathematical toy. The paradigm asserting that the universe in not locally real, implies that $p$-adic QM allows realism.
Quantum computing and the non-locality tests can be formulated using $p$-adic QM. This idea is already presented in [21]. Here, we give a quick review. The standard formulation of these theories uses QM on the Hilbert space $\mathbb{C}^{N}$. We select a prime number $p$ and a positive integer $l$, and a subset $G_{l}^{0}\subseteq G_{l}$, with cardinality $G_{l}^{0}=N$. Notice that we require that $N\leq p^{l}$. Now, the set
(3.1) $\left\{p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-I\right|_{p}\right);I\in G_{l}^{0}\right\}\subset\mathcal{D}_{l}(\mathbb{Z}_{p})$
is orthonormal set with respect to the inner product
$\left\langle\varphi,\phi\right\rangle={\displaystyle\int\limits_{\mathbb{Z}_{p}}}\varphi\left(x\right)\overline{\phi}\left(x\right)dx.$
We denote by $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ the $N$-dimensional Hilbert space spanned by (3.1). Then, $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ is isomorphic to $\mathbb{C}^{N}$, as Hilbert spaces. Notice that the isomorphism exists for any $p$ and $l$ such that $N\leq p^{l}$.
The advantage of using $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ over $\mathbb{C}^{N}$ is that we can construct, in a natural way, continuos versions of quantum models in $\mathbb{C}^{N}$. We identify $p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ with $e_{I}$, where $\left\{e_{I}\right\}_{I\in G_{l}^{0}}$ is the canonical basis of $\mathbb{C}^{N}$. Let
$\left[H_{J,K}\right]_{1\leq J,K\leq N}=\left[H_{J,K}\right]_{J,K\in G_{l}^{0}}$
be a Hermitian matrix. We attach to this matrix the kernel
$h(x,y):=p^{l}{\displaystyle\sum\limits_{J\in G_{l}^{0}}}{\displaystyle\sum\limits_{K\in G_{l}^{0}}}H_{J,K}\Omega\left(p^{l}\left|x-J\right|_{p}\right)\Omega\left(p^{l}\left|y-K\right|_{p}\right),$
for $x,y\in\mathbb{Z}_{p}$, and the linear operator
(3.2) $\varphi\left(x\right)\rightarrow\boldsymbol{H}\varphi\left(x\right)={\displaystyle\int\limits_{\mathbb{Z}_{p}}}h(x,y)\varphi\left(y\right)dy,$
where $\varphi\left(y\right)=\sum_{K\in G_{l}^{0}}c_{K}\Omega\left(p^{l}\left|y-K\right|_{p}\right)\in\mathcal{\chi}_{N}(\mathbb{Z}_{p})$.  Now, if we take $\varphi\in\mathcal{C}(\mathbb{Z}_{p})$ to be a continuous function defined on $\mathbb{Z}_{p}$, then, the integral in (3.2) gives a function from $\mathcal{C}(\mathbb{Z}_{p})$. Furthermore, $\mathcal{C}(\mathbb{Z}_{p})$ is dense in $L^{2}(\mathbb{Z}_{p})$, and
$\left\|\boldsymbol{H}\varphi\right\|_{L^{2}(\mathbb{Z}_{p})}\leq\left\|h\right\|_{L^{2}(\mathbb{Z}_{p}\times\mathbb{Z}_{p})}\left\|\varphi\right\|_{L^{2}(\mathbb{Z}_{p})}.$
Therefore, $\boldsymbol{H}$ extends  to a linear, bounded operator on $L^{2}(\mathbb{Z}_{p})$. Notice that this operator is non-local, and that by construction $\boldsymbol{H}:\mathcal{\chi}_{N}(\mathbb{Z}_{p})\rightarrow\mathcal{\chi}_{N}(\mathbb{Z}_{p})$; this operator is represented by the matrix $\left[H_{J,K}\right]_{1\leq J,K\leq N}$. It is not difficult to verify that $\boldsymbol{H}$ is symmetric, and consequently self-adjoint. In conclusion,
$i\frac{\partial}{\partial t}\Psi\left(x,t\right)=\boldsymbol{H}\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p}\text{, }t\geq 0,$
is a continuous Schrödinger equation attached to matrix $\left[H_{J,K}\right]_{1\leq J,K\leq N}$. For further details, the reader may consult [21]; see also [24], [25], and the references therein.
In 1964, Bell established that quantum mechanics is inherently non-local through a non-locality test known now as Bell inequalities, [2]. Later, non-locality tests without using inequalities were established by several authors; see, e.g., [3], [30]. A relevant observation is that QM on $L^{2}(\mathcal{\chi}_{N}(\mathbb{Z}_{p}))\subset L^{2}(\mathbb{Z}_{p})$ is a non-local theory, which naturally agrees with Bell’s conclusion.
4. QM is a non-local theory that admits realism
[…]
4.1. An orthonormal basis for $L^{2}(\mathbb{Q}_{p})$
Any $p$-adic number $x\neq 0$ has a unique expansion of the form $x=p^{ord(x)}\sum_{j=0}^{\infty}x_{j}p^{j},$ where $x_{j}\in\{0,\dots,p-1\}$ and $x_{0}\neq 0$. By using this expansion, we define the fractional part of $x\in\mathbb{Q}_{p}$, denoted $\{x\}_{p}$, as the rational number
$\left\{x\right\}_{p}=\left\{\begin{array}[c]{lll}0&\text{if}&x=0\text{ or }ord(x)\geq 0\\ &&\\ p^{ord(x)}\sum_{j=0}^{-ord(x)-1}x_{j}p^{j}&\text{if}&ord(x)<0.\end{array}\right$.
Set $\chi_{p}(y)=\exp(2\pi i\{y\}_{p})$ for $y\in\mathbb{Q}_{p}$. The map $\chi_{p}(\cdot)$ is an additive character on $\mathbb{Q}_{p}$, i.e., a continuous map from $\left(\mathbb{Q}_{p},+\right)$ into $S$ (the unit circle considered as a multiplicative group) satisfying $\chi_{p}(x_{0}+x_{1})=\chi_{p}(x_{0})\chi_{p}(x_{1})$, $x_{0},x_{1}\in\mathbb{Q}_{p}$; see, e.g., [42, Section 2.3].
We set
$\mathbb{Q}_{p}/\mathbb{Z}_{p}=\left\{\sum_{j=-1}^{-m}x_{j}p^{j};\text{for some }m>0\right\}.$
For $b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$, $r\in\mathbb{Z}$, we denote by $\Omega\left(\left|p^{r}x_{p}-b\right|_{p}\right)$ the characteristic function of the ball $bp^{-r}+p^{-r}\mathbb{Z}_{p}$.
We now define
$\psi_{rbk}\left(x_{p}\right)=p^{\frac{-r}{2}}\chi_{p}(p^{-1}k\left(p^{r}x_{p}-b\right))\Omega\left(\left|p^{r}x_{p}-b\right|_{p}\right),$
where $r\in\mathbb{Z}$, $k\in\{1,\dots,p-1\}$, and $b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$.
Then, $\left\{\psi_{rbk}\left(x_{p}\right)\right\}_{rbk}$ forms a complete orthonormal basis of $L^{2}(\mathbb{Q}_{p})$, and
$\boldsymbol{D}^{\alpha}\psi_{rbk}\left(x\right)=p^{\left(1-r\right)\alpha}\psi_{rbk}\left(x\right)\text{, for any }r,b,k;$
see, e.g., [42, Theorems 9.4.5 and 8.9.3], or [44, Theorem 3.3]
4.2. Some results about tensor product of Hilbert spaces
We set $L^{2}(\mathbb{R},dx_{\infty})=L^{2}(\mathbb{R})$, where  denotes the Lebesgue measure, and $L^{2}\left(\mathbb{Q}_{p},dx_{p}\right)=L^{2}\left(\mathbb{Q}_{p}\right)$, where $dx_{p}$ denotes the Haar measure of $\mathbb{Q}_{p}$. We denote by $dx_{\infty}dx_{p}$, the product measure of $dx_{\infty}$ and $dx_{p}$ on $\mathbb{R}\times\mathbb{Q}_{p}$. On the other hand, $L^{2}\left(\mathbb{Q}_{p}\right)$ is a separable Hilbert space because it has a countable orthonormal basis $\left\{\psi_{rbk}\left(x_{p}\right)\right\}_{rbk}$, and also $L^{2}(\mathbb{R})$ is a separable Hilbert space. The Haar wavelets $\left\{H_{m}(x_{\infty})\right\}_{m\in\mathbb{N}}$, which are functions with compact support, form an orthonormal basis of $L^{2}(\mathbb{R})$; also the Hermite polynomials, which are Schwartz functions in $\mathbb{R}$, form an orthonormal basis of $L^{2}(\mathbb{R})$. Then $\left\{H_{m}\left(x_{\infty}\right)\psi_{rbk}\left(x_{p}\right)\right\}_{rbk,m}$ is an orthonormal basis of $L^{2}(\mathbb{R})\otimes L^{2}\left(\mathbb{Q}_{p}\right)$, cf. [54, Proposition 2.7] or [55, Section II.4, Proposition 2], where $\otimes$ denotes the tensor product of Hilbert spaces.
There is a unique isomorphism from $L^{2}(\mathbb{R},dx_{\infty})\otimes L^{2}\left(\mathbb{Q}_{p},dx_{p}\right)$ to $L^{2}\left(\mathbb{R}\times\mathbb{Q}_{p},dx_{\infty}dx_{p}\right)$ such that $f{\textstyle\bigotimes}g$ is taken into $fg$; and there is a unique isomorphism from $L^{2}\left(\mathbb{R}\times\mathbb{Q}_{p},dx_{\infty}dx_{p}\right)$ to $L^{2}\left(\mathbb{R},dx_{\infty};L^{2}\left(\mathbb{Q}_{p},dx_{p}\right)\right)$ such that $f(x_{\infty},x_{p})$ is taken into the function $x_{p}\rightarrow f(\cdot,x_{p})$. Similarly, there is a unique isomorphism from $L^{2}\left(\mathbb{R}\times\mathbb{Q}_{p},dx_{\infty}dx_{p}\right)$ to $L^{2}\left(\mathbb{Q}_{p},dx_{p};L^{2}\left(\mathbb{R},dx_{\infty}\right)\right)$ such that $f(x_{\infty},x_{p})$ is taken into the function $x_{\infty}\rightarrow f(x_{\infty},\cdot)$; see [55, Section II.4, Theorem II.10].
[…]
Take $\Psi\left(t\right)=\Psi\left(x_{\infty},x_{p},t\right)\in L^{2}(\mathbb{R})\otimes L^{2}\left(\mathbb{Q}_{p}\right)$. For a fixed value of $x_{\infty}$, we set $\Psi\left(x_{\infty},x_{p},t\right)=\Psi_{p}\left(x_{p},t\right)$; and for a fixed value of $x_{p}$, we set $\Psi\left(x_{\infty},x_{p},t\right)=\Psi_{\infty}\left(x_{\infty},t\right)$. These functions satisfy $\Psi_{\infty}\left(\cdot,t\right)\in L^{2}(\mathbb{R})$, $\Psi_{p}\left(\cdot,t\right)\in L^{2}\left(\mathbb{Q}_{p}\right)$ for any $t\in\mathbb{R}$ .
Now, given a wavefunction $\Psi\left(x_{\infty},x_{p},t\right)$, the Born rule produces several different probability distributions.
Type 1
We fix $x_{p}$, then
$\frac{\left|\Psi\left(x_{\infty},x_{p},t\right)\right|^{2}}{\left\|\Psi\left(\cdot,x_{p},t\right)\right\|_{L^{2}(\mathbb{R})}^{2}}dx_{\infty}=\frac{\Psi_{\infty}\left(x_{\infty},t\right)}{\left\|\Psi_{\infty}\left(\cdot,t\right)\right\|_{L^{2}(\mathbb{R})}^{2}}dx_{\infty}$
is a probability measure in $\left(\mathbb{R},\mathcal{B}(\mathbb{R})\right)$, for $t\in\mathbb{R}$, where $\mathcal{B}(\mathbb{R})$ is the Borel $\sigma$-algebra of $\mathbb{R}$.
Type 2
We fix $x_{\infty}$, then
$\frac{\left|\Psi\left(x_{\infty},x_{p},t\right)\right|^{2}}{\left\|\Psi\left(x_{\infty},\cdot,t\right)\right\|_{L^{2}(\mathbb{Q}_{p})}^{2}}dx_{p}=\frac{\Psi_{p}\left(x_{p},t\right)}{\left\|\Psi_{p}\left(\cdot,t\right)\right\|_{L^{2}(\mathbb{Q}_{p})}^{2}}dx_{p}$
is a probability measure in $\left(\mathbb{Q}_{p},\mathcal{B}(\mathbb{Q}_{p})\right)$, for $t\in\mathbb{R}$, where $\mathcal{B}(\mathbb{Q}_{p})$ is the Borel $\sigma$-algebra of $\mathbb{Q}_{p}$.
Type 3
For $t\in\mathbb{R}$,
[…]
There are three different quantum states, $\Psi\left(x_{\infty},x_{p},t\right)$, $\Psi_{\infty}\left(x_{\infty},t\right)$, $\Psi_{p}\left(x_{p},t\right)$; we argue that they provide, in principle, three different pictures of the quantum reality. Notice, that in general, $\Psi\left(x_{\infty},x_{p},t\right)\neq\Psi_{\infty}\left(x_{\infty},t\right)\Psi_{p}\left(x_{p},t\right)$. We will come back to this discussion in Section 7.
4.4. Non-local Schrödinger equations
The next step is to determine the unitary evolution operators acting on $L^{2}(\mathbb{R})\otimes L^{2}\left(\mathbb{Q}_{p}\right)$, which may have a physical meaning. We assume that the standard QM on $L^{2}\left(\mathbb{R}^{N}\right)$ provides a good description of quantum phenomena, when $\mathbb{R}\times\mathbb{R}^{N}$ is a good model for the physical configuration space associated with the phenomena. We also assume that $p$-adic QM on $L^{2}\left(\mathbb{Q}_{p}^{N}\right)$ accurately describes quantum phenomena, when $\mathbb{R}\times\mathbb{Q}_{p}^{N}$ is a good model for the physical configuration space. By using a model of the form $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{Q}_{p}\right)^{N}$, we have regions of type
$\mathbb{R}\times\left(\mathbb{R}\times\left\{\alpha\right\}\right)^{N}\simeq\mathbb{R}\times\mathbb{R}^{N},$
for some fixed $\alpha\in\mathbb{Q}_{p}$, where the standard (local) QM is valid, and regions of the form
$\mathbb{R}\times\left(\left\{\beta\right\}\times\mathbb{Q}_{p}\right)^{N}\simeq\mathbb{R}\times\mathbb{Q}_{p}^{3N},$
for some fixed $\beta\in\mathbb{R}$, where $p$-adic (non-local) QM is valid.
[…]
$\left\{\begin{array}[c]{l}i\frac{\partial}{\partial t}\Psi_{\infty}\left(x_{\infty},t\right)=\boldsymbol{H}_{\infty}\Psi_{\infty}\left(x_{\infty},t\right)\text{, \ }x_{\infty}\in\mathbb{R},t\geq 0;\\ \\ \Psi_{\infty}\left(\cdot,t\right)\in L^{2}\left(\mathbb{R}\right),\text{ }t\geq 0;\\ \\ \left\|\Psi_{p}\left(\cdot,t\right)\right\|_{2}=1,\text{ }t\geq 0,\end{array}\right$.
and
$\left\{\begin{array}[c]{l}i\frac{\partial}{\partial t}\Psi_{p}\left(x_{p},t\right)=\boldsymbol{H}_{p}\Psi_{p}\left(x_{p},t\right)\text{, \ }x_{p}\in\mathbb{Q}_{p},t\geq 0;\\ \\ \Psi_{p}\left(\cdot,t\right)\in L^{2}\left(\mathbb{Q}_{p}\right),\text{ }t\geq 0;\\ \\ \left\|\Psi_{p}\left(\cdot,t\right)\right\|_{2}=1,\text{ }t\geq 0,\end{array}\right$.
where $\boldsymbol{H}_{p}:L^{2}\left(\mathbb{Q}_{p}\right)\supset Dom(\boldsymbol{H}_{\infty})\rightarrow L^{2}\left(\mathbb{Q}_{p}\right)$, and $\boldsymbol{H}_{\infty}:L^{2}\left(\mathbb{R}\right)\supset Dom(\boldsymbol{H}_{\infty})\rightarrow L^{2}\left(\mathbb{R}\right)$ are Hamiltonians, and $\boldsymbol{H}_{p}$ is a non-local operator, while $\boldsymbol{H}_{\infty}$ is a local operator.
A straightforward calculation shows that $\Psi\left(x_{\infty},x_{p},t\right)$ obeys to the following Schrödinger equation:
[…]
where the Hamiltonian $\boldsymbol{H}_{\infty}+\boldsymbol{H}_{p}:L^{2}\left(\mathbb{R}\right){\textstyle\bigotimes}L^{2}\left(\mathbb{Q}_{p}\right)\supset Dom(\boldsymbol{H}_{\infty}+\boldsymbol{H}_{p})\rightarrow L^{2}\left(\mathbb{R}\right){\textstyle\bigotimes}L^{2}\left(\mathbb{Q}_{p}\right)$ is a non-local operator.
Non-local operators of type $\boldsymbol{H}_{\infty}+\boldsymbol{H}_{p}$ (of course with $\boldsymbol{H}_{p}\neq 0$) allow spooky action at a distance. Then, QM (in the sense of Dirac-von Neumann) on the Hilbert space $L^{2}\left(\mathbb{R}\right){\textstyle\bigotimes}L^{2}\left(\mathbb{Q}_{p}\right)$, with time evolution $\left\{e^{-it\left(\boldsymbol{H}_{\infty}+\boldsymbol{H}_{p}\right)}\right\}_{t\in\mathbb{R}}$ is a non-local theory; consequently, the fact the universe is not locally real, implies that this QM admits realism.
Remark 1.
The above argument only gives a subset of possible admissible Hamiltonians for our quantum formalism. Notice that any operator of the form
$\boldsymbol{H}=\boldsymbol{H}_{\infty}+\boldsymbol{P}\left(\boldsymbol{H}_{p},\boldsymbol{H}_{\infty}\right)+\boldsymbol{H}_{p},$
where operator $\boldsymbol{P}$ satisfies
$\boldsymbol{P}\left(\boldsymbol{0},\boldsymbol{H}_{\infty}\right)=\boldsymbol{P}\left(\boldsymbol{H}_{p},\boldsymbol{0}\right)=\boldsymbol{0}$
should be considered as admissible.
We now assume that $\left\{\psi_{rbk}\left(x_{p}\right)\right\}_{rbk}$ is an orthonormal basis of $L^{2}\left(\mathbb{Q}_{p}\right)$ consisting of eigenfunctions of $\boldsymbol{H}_{p}$,
$\boldsymbol{H}_{p}\psi_{rbk}=E_{rbk}\psi_{rbk}\text{, with }E_{rbk}\in\mathbb{R}\text{,}$
and that $\left\{\theta_{m}\left(x_{\infty}\right)\right\}_{m}$ is an orthonormal basis of $L^{2}\left(\mathbb{R}\right)$ consisting of eigenfunctions of $\boldsymbol{H}_{\infty}$,
$\boldsymbol{H}_{\infty}\theta_{m}=E_{m}\theta_{m}\text{, with }E_{m}\in\mathbb{R}\text{.}$
Then, any solution of the Schrödinger equation in  (4.1) has the form
$\Psi\left(x_{\infty},x_{p},t\right)={\displaystyle\sum\limits_{rbk}}{\displaystyle\sum\limits_{m}}A_{rbk,m}e^{-it\left(E_{rbk}+E_{m}\right)}\theta_{m}\left(x_{\infty}\right)\psi_{rbk}\left(x_{p}\right).$
Since $\left\{\theta_{m}\left(x_{\infty}\right)\psi_{rbk}\left(x_{p}\right)\right\}_{rbk,m}$ is an orthonormal basis of $L^{2}(\mathbb{R})\otimes L^{2}\left(\mathbb{Q}_{p}\right)\simeq L^{2}(\mathbb{R}\otimes\mathbb{Q}_{p})$, $\Psi\left(x_{\infty},x_{p},t\right)\in L^{2}(\mathbb{R}\otimes\mathbb{Q}_{p})$, for any $t\geq 0$. We now fix the coefficients $A_{rbk,m}$ so that
$\Psi\left(x_{\infty},x_{p},0\right)={\displaystyle\sum\limits_{rbk}}{\displaystyle\sum\limits_{m}}A_{rbk,m}\theta_{m}\left(x_{\infty}\right)\psi_{rbk}\left(x_{p}\right).$
Finally, using that $\left\{\theta_{m}\left(x_{\infty}\right)\psi_{rbk}\left(x_{p}\right)\right\}_{rbk,m}$ is an orthonormal basis of $L^{2}(\mathbb{R}\otimes\mathbb{Q}_{p})$,
$\left\|\Psi\left(x_{\infty},x_{p},t\right)\right\|_{L^{2}(\mathbb{R}\otimes\mathbb{Q}_{p})}^{2}={\displaystyle\sum\limits_{rbk}}{\displaystyle\sum\limits_{m}}\left|A_{rbk,m}\right|^{2}=\left\|\Psi\left(x_{\infty},x_{p},0\right)\right\|_{L^{2}(\mathbb{R}\otimes\mathbb{Q}_{p})}^{2}=1\text{,}$
for $t\geq 0$.
5. The measurement problem and GWR theory
In this section, we review quickly the measurement problem and the Ghirardi-Rimini-Weber (GRW) collapse theory. In the next section, we propose a solution to the measurement problem; the collapse mechanism resembles the one in the GRW theory, but there are several fundamental differences.
5.1. The measurement problem
We review the measurement problem following Bell, [31], and Landau and Lifshitz, [32].
We consider a system consisting of two parts: a classical apparatus (a macroscopic system) and a quantum object (a microscopic system). The process of measurement involves the interaction of these two parts; as a result the apparatus passes from its initial state into some other. From this change of state one draws conclusions concerning the state of the quantum object. Let $g$ be the pointer of the apparatus $\mathcal{A}$, i.e., the readings of it. Let $\widehat{\boldsymbol{g}}$ be the self-adjoint operator corresponding to the observable $g$, which we suppose to have a discrete and non-degenerate spectrum. Let $g_{n}\in\mathbb{R}$, $F_{n}\in\mathcal{H}_{\mathcal{A}}$, $n=1,2,\ldots$, be, respectively, the eigenvalues and eigenfunctions of $\widehat{\boldsymbol{g}}$, where $\mathcal{H}_{\mathcal{A}}$ is the state space of $\mathcal{A}$. We emphasize that the wavefunctions of the apparatus $F_{n}=F_{n}\left(\xi,t\right)$, $\left(\xi,t\right)\in\mathbb{R}^{N}\times\mathbb{R}$ are functions in the configuration space $\mathbb{R}^{N}\times\mathbb{R}$, and thus by our previous discussion the apparatus is in the macroscopic realm. Let $F_{0}\left(\xi\right)=F_{n_{0}}\left(\xi,0\right)$ be the wavefunction of the initial state of the apparatus at $t=0$ before the measurement.
Let $\psi_{n}$, $n=1,2,\ldots$ , be the eigenvectors of the operator $\widehat{\boldsymbol{o}}$ corresponding to the observable $o$ that is to be determined by the apparatus. We assume that $\psi_{n}=\psi_{n}\left(q,t\right)$, $\left(q,t\right)\in\mathcal{X}^{N}\times\mathbb{R}$,  meaning that each $\psi_{n}\left(q,t\right)$ is the state of a quantum object in $\mathcal{X}^{N}\times\mathbb{R}$.  Only in this section, $\mathcal{X}$ is an arbitrary space, so it can be $\mathbb{R}$ or $\mathbb{Q}_{p}$.
Let
$\Psi_{0}\left(q\right)={\displaystyle\sum\limits_{n}}c_{n}\psi_{n}\left(q,0\right)$
be the initial state of the quantum object. The initial state of the whole system before the measurement is
$\Psi_{0}{\textstyle\bigotimes}F_{0}=\Psi_{0}\left(q\right)F_{0}\left(\xi\right)\in\mathcal{H}_{\mathcal{S}}{\textstyle\bigotimes}\mathcal{H}_{\mathcal{A}},$
where $\mathcal{H}_{\mathcal{S}}$ is the Hilbert space corresponding to the quantum object $\mathcal{S}$. Due to the interaction between $\mathcal{S}$ and $\mathcal{A}$, this wavefunction evolves according the Schrödinger equation. We denote by $\Psi\left(q,\xi,t\right)$ the wavefunction of the system $\mathcal{S}+\mathcal{A}$ at time $t$, then
(5.1) $\Psi\left(q,\xi,t\right)={\displaystyle\sum\limits_{n}}a_{n}\psi_{n}\left(q,t\right)F_{n}\left(\xi,t\right)={\displaystyle\sum\limits_{n}}A_{n}\left(q,t\right)F_{n}\left(\xi,t\right).$
The essence of the measurement problem is that after the interaction of $\mathcal{S}$, and $\mathcal{A}$, the Schrödinger equation does not give a definite post-interaction position for the measuring apparatus pointer. In QM, this problem is fixed by introducing the “collapse postulate”: when the measurement occurs, the wavefunction ceases to evolve according to the Schrödinger equation momentarily does something entirely different.
Following Landau and Lifshitz, [32], the classical nature of the apparatus implies that at any instant $t>0$, the quantity $g$ (the reading of the apparatus) has a definite value, and thus after the measurement the system $\mathcal{S}+\mathcal{A}$ will be not described by (5.1), but by the term corresponding to the reading $g_{n}$ of the apparatus,
$A_{n}\left(q,t\right)F_{n}\left(\xi,t\right).$
By writing
$A_{n}\left(q,t\right)=\left\|A_{n}\left(\cdot,t\right)\right\|_{2}\frac{A_{n}\left(q,t\right)}{\left\|A_{n}\left(\cdot,t\right)\right\|_{2}},$
one gets that $\frac{A_{n}\left(q,t\right)}{\left\|A_{n}\left(\cdot,t\right)\right\|_{2}}$ is the normalized wavefunction of $\mathcal{S}$ after the measurement.
5.2. The GRW collapse model
Ghirardi-Rimini-Weber (GRW) theory, and other spontaneous collapse theories, solve the quantum measurement problem by modifying the Schrödinger equation with nonlinear and stochastic terms that cause the wavefunction to collapse spontaneously. We give a quick review of the GRW collapse model; see, e.g., [33]-[36], and the references therein. This section is based on [33] and [36].
In the GRW proposal, the wavefunctions evolve according to the Schrödinger equation most of the time, except for occasional random moments when they suffer a spontaneous localization. Consider the Gaussian function
$g_{r}\left(x\right)=\frac{1}{\left(2\pi\sigma^{2}\right)^{\frac{1}{4}}}e^{-\frac{\left(x-r\right)^{2}}{4\sigma^{2}}}\text{, with }{\displaystyle\int\limits_{\mathbb{R}}}g_{r}^{2}\left(x\right)dx=1,$
centered at the point $x=r$. During the spontaneous localization at time $t$, the wavefunction gets multiplied by $g_{r}\left(x\right)$:
(5.2) $\Psi\left(x,t_{+}\right)=\frac{g_{r}\left(x\right)\Psi\left(x,t_{-}\right)}{P(r)},$
where $\Psi\left(x,t_{-}\right)$ is the wavefunction before time $t$, $\Psi\left(x,t_{+}\right)$ is the wavefunction after time $t$, and
$P(r)=\sqrt{{\displaystyle\int\limits_{\mathbb{R}}}\left|g_{r}\left(x\right)\Psi\left(x,t_{-}\right)\right|^{2}dx}.$
The parameter $r$ is random with a probability distribution $P(r)$. The localization (5.2) occurs at random times, with a Poisson distribution with mean frequency $\lambda$. The parameters $\lambda$ and $\sigma$ are phenomenological constants that should be understood as new constants of Nature. The theory can be formulated through a short list of principles, from these using the statistical operator formalism, a master equation for $N$ particles can be obtained. This equation is a substitute to the standard Schrödinger equation; see [35], and the references therein.
It is relevant that $\Psi\left(x,t_{-}\right)\rightarrow\Psi\left(x,t_{+}\right)$ tends to localize $\Psi\left(x,t_{-}\right)$ because the Gaussian function is concentrated around $x=r$ for a suitable choice of $\sigma$. We invite the reader to review the several examples presented in Norsen’s book [33, Chapter 9]. The collapse mechanism in the GWR theory resembles the one introduced here. However, the perspective and the implications of the GWR theory are radically different from those of our theory, as we shall discuss in the coming sections.
6. The collapse (localization) of the wavefunctions
We now apply the quantum mechanical formalism developed in the previous sections to propose a solution to the collapse of the wavefunction. In our formalism, the collapse of the wavefunction is a consequence of the geometry of the space-time (the configuration space).
We use the notation introduced at the end of Section 4. The apparatus is in an space-time that can be modeled as $\mathbb{R}\times\mathbb{R}$, where the locality and realism are valid; while the quantum particle is in an space-time that can be modeled as $\mathbb{R}\times\mathbb{Q}_{p}$, where non-locality and realism are valid. The wavefunctions of the quantum particle are of the form $\Psi_{p}\left(x_{p},t\right)\in L^{2}\left(\mathbb{Q}_{p}\right)$, for any $t\geq 0$, and
(6.1) $\Psi_{p}\left(x_{p},t\right)={\displaystyle\sum\limits_{rbk}}c_{rbk}\left(t\right)\psi_{rbk}\left(x_{p}\right);$
while the wavefunctions of the apparatus are of the form $\psi_{\infty}\left(x_{\infty},t\right)\in L^{2}\left(\mathbb{R}\right)$, for any $t\geq 0$, and
(6.2) $\Psi_{\infty}\left(x_{\infty},t\right)={\displaystyle\sum\limits_{m}}c_{m}\left(t\right)\theta_{m}\left(x_{\infty}\right).$
The wavefunction of the system $\mathcal{S}+\mathcal{A}$ at time $t$, is given by
(6.3) $\Psi\left(x_{\infty},x_{p},t\right)=\Psi_{\infty}\left(x_{\infty},t\right)\Psi_{p}\left(x_{p},t\right)={\displaystyle\sum\limits_{rbk}}{\displaystyle\sum\limits_{m}}a_{rbk,m}\left(t\right)\psi_{rbk}\left(x_{p}\right)\theta_{m}\left(x_{\infty}\right),$
where $a_{rbk,m}\in\mathbb{C}$. Then, (6.3) is exactly (5.1), when we use the space-time model $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{Q}_{p}\right)$. The time evolution of the quantum state $\Psi\left(x_{\infty},x_{p},t\right)$ obeys to a non-local Schrödinger equation on $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{Q}_{p}\right)$.
6.1. Key calculations
The Monna map is defined as
$\begin{array}[c]{cccc}\mathcal{M}:&\mathbb{Q}_{p}&\rightarrow&\mathbb{R}_{\geq 0}\\ &&&\\ &x_{p}={\displaystyle\sum\limits_{j=\gamma}^{\infty}}y_{j}p^{j}&\rightarrow&x_{\infty}={\displaystyle\sum\limits_{j=\gamma}^{\infty}}y_{j}p^{-j-1}.\end{array}$
This map captures the strangeness of QM: $\mathbb{R}$ does not contain a copy of $\mathbb{Q}_{p}$ that preserves both the topology and algebraic structure of $\mathbb{Q}_{p}$. More precisely, $\mathbb{R}$ is not isomorphic to $\mathbb{Q}_{p}$ as local fields, [57, Chpater I, Section 3]. The Monna map is a continuous, surjective, but not injective, [42, Section 1.9.4].
At the measurement time, the apparatus scan a compact region in the microscopic realm ($\mathbb{Q}_{p}$), the description of the classical interaction between the apparatus and the quantum particle requires the identification of points of $\mathbb{Q}_{p}$ with points in $\mathbb{R}$. A such identification is obtained through the Monna map.
We denote by $dx_{\infty}$, the Lebesgue measure on $\left(\mathbb{R},+\right)$ and by $dx_{p}$, the Haar measure of $\left(\mathbb{Q}_{p},+\right)$. The mapping $\mathcal{M}$ is measurable, because it is continuous, i.e.,
$\mathcal{M}:\left(\mathbb{Q}_{p},\mathcal{B}\left(\mathbb{Q}_{p}\right)\right)\rightarrow\left(\mathbb{R}_{\geq 0},\mathcal{B}\left(\mathbb{R}_{\geq 0}\right)\right),$
and thus for any $A\in\mathcal{B}\left(\mathbb{R}_{\geq 0}\right)$,
${\displaystyle\int\limits_{A}}dx_{\infty}={\displaystyle\int\limits_{\mathcal{M}^{-1}\left(A\right)}}dx_{p},$
[…]
(6.4) ${\displaystyle\int\limits_{A}}F\left(x_{\infty}\right)dx_{\infty}={\displaystyle\int\limits_{\mathcal{M}^{-1}\left(A\right)}}F\left(\mathcal{M}\left(x_{p}\right)\right)dx_{p}.$
6.1.1. Formula 1
For each $t\in\mathbb{R}$,
$\frac{\left|\Psi_{p}\left(x_{p},t\right)\right|^{2}\left|\Psi_{\infty}\left(\mathcal{M}\left(x_{p}\right),t\right)\right|^{2}}{\left\|\Psi_{p}\left(\mathcal{\cdot},t\right)\Psi_{\infty}\left(\mathcal{M}\left(\cdot\right),t\right)\right\|_{L^{2}(\mathbb{Q}_{p})}^{2}}dx_{\infty}=\frac{1}{A\left(\mathcal{M}\right)}\left|\Psi_{p}\left(x_{p},t\right)\right|^{2}\left|\Psi_{\infty}\left(\mathcal{M}\left(x_{p}\right),t\right)\right|^{2}dx_{p}$
is a probability measure in $\left(\mathbb{Q}_{p},\mathcal{B}(\mathbb{Q}_{p})\right)$, where $\mathcal{B}(\mathbb{Q}_{p})$ is the Borel $\sigma$-algebra of $\mathbb{Q}_{p}$. More precisely, for a Borel subset $B\subset\mathbb{Q}_{p}$, its probability measure is given by
(6.5) $P_{\text{int}}(B,t)=\frac{1}{A\left(\mathcal{M}\right)}{\displaystyle\int\limits_{B}}\text{ \ }\left|\Psi_{p}\left(x_{p},t\right)\right|^{2}\left|\Psi_{\infty}\left(\mathcal{M}\left(x_{p}\right),t\right)\right|^{2}dx_{p}.$
$P_{\text{int}}(B,t)$ is the probability of interaction of the quantum states $\Psi_{\infty}\left(x_{\infty},t\right)$ and $\Psi_{p}\left(x_{p},t\right)$ at the time $t$ in the region $B\subset\mathbb{Q}_{p}$. If $\Psi_{\infty}\left(x_{\infty},t\right)$ is the state of of the apparatus and $\Psi_{p}\left(x_{p},t\right)$ is the state of the quantum object at the time $t$, then $P_{\text{int}}(B,t)$ gives the probability of interaction between the apparatus and the quantum object, when the apparatus scans a region $B\subset\mathbb{Q}_{p}$ at the the time $t$. The formula for $P_{\text{int}}(B,t)$ follows from (6.4).
6.1.2. Formula 2
We assume that the scanned region $B$ is a ball, $B=p^{l}a+p^{l}\mathbb{Z}_{p}$, for $l\in\mathbb{N}$, $a\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. Then
$\mathcal{M}\left(B\right)=\mathcal{M}(p^{l}a)+\left[0,p^{-l}\right]$
is an interval in $\mathbb{R}_{\geq 0}$; see [42, Lemma 1.9.11].
We now use expansion (6.1), where the support of $\left\{\psi_{rbk}\right\}_{rbk}$ (denoted as supp $\psi_{rbk}$) is a ball. We now recall that in $p$-adic topology two balls are disjoint or one in contained in the other; consequently, there are two cases:
[…]
or
(6.7) $\mathrm{supp}\text{ }\psi_{rbk}\cap B=\varnothing\Leftrightarrow\mathrm{supp}\text{ }\psi_{rbk}\subset B^{c},$
where $B^{c}=\mathbb{Q}_{p}\smallsetminus B$. In case (6.6), $\mathrm{supp}$ $\psi_{rbk}\cap B\neq\varnothing$. Using this fact, we rewrite (6.1) as
(6.8) $\Psi_{p}\left(x_{p},t\right)=\Psi_{p}^{\left(B\right)}\left(x_{p},t\right)+\Psi_{p}^{\left(B^{c}\right)}\left(x_{p},t\right),$
where
(6.9) $\Psi_{p}^{\left(B\right)}\left(x_{p},t\right):={\displaystyle\sum\limits_{\begin{subarray}{c}rbk\\ \mathrm{supp}\text{ }\psi_{rbk}\cap B\neq\varnothing\end{subarray}}}c_{rbk}(t)\psi_{rbk}\left(x_{p}\right)\in\mathcal{H}_{\mathcal{S}},$
and
(6.10) $\Psi_{p}^{\left(B^{c}\right)}\left(x_{p},t\right)={\displaystyle\sum\limits_{\begin{subarray}{c}rbk\\ \mathrm{supp}\text{ }\psi_{rbk}\cap B=\varnothing\end{subarray}}}c_{rbk}(t)\psi_{rbk}\left(x_{p}\right)\in\mathcal{H}_{\mathcal{S}}.$
Furthermore, $\left\langle\Psi_{p}^{\left(B\right)}\left(x_{p},t\right),\Psi_{p}^{\left(B^{c}\right)}\left(x_{p},t\right)\text{ }\right\rangle=0$, for any $t\geq 0$, and
$1_{B}\left(x_{p}\right)\Psi_{p}\left(x_{p},t\right)=\Psi_{p}^{\left(B\right)}\left(x_{p},t\right).$
Therefore,
${\displaystyle\int\limits_{B}}\left|\Psi_{p}\left(x_{p},t\right)\right|^{2}dx_{p}={\displaystyle\int\limits_{\mathbb{Q}_{p}}}\left|\Psi_{p}^{\left(B\right)}\left(x_{p},t\right)\right|^{2}dx_{p}.$
In conclusion, the multiplication by $1_{B}\left(x_{p}\right)$ induces the map
(6.11) $\Psi_{p}\left(x_{p},t\right)=\Psi_{p}^{\left(B\right)}\left(x_{p},t\right)+\Psi_{p}^{\left(B^{c}\right)}\left(x_{p},t\right)\rightarrow\Psi_{p}^{\left(B\right)}\left(x_{p},t\right),$
where $\Psi_{p}^{\left(B\right)}\left(x_{p},t\right)$ satisfies (6.11). During the scanning process, the apparatus localized the wave function, $\Psi_{p}\left(x_{p},t\right)$, in the region $B\subset\mathbb{Q}_{p}$. It is essential that the localization process extract exactly the part of the wavefunction that describes the probability density for the particle to be in $B$ at any time.
6.1.3. Formula 3
We now compute the restriction of $\psi_{rbk}\left(x_{p}\right)$ to the ball $B=p^{l}a+p^{l}\mathbb{Z}_{p}$:
(6.12) $\Omega\left(p^{l}\left|x_{p}-p^{l}a\right|_{p}\right)\psi_{rbk}\left(x_{p}\right)=\left\{\begin{array}[c]{ll}\psi_{rbk}\left(x_{p}\right)&\text{if }bp^{-r}-p^{l}a\in p^{l}\mathbb{Z}_{p}\text{, }r\leq-l\\ &\\ p^{\frac{-r}{2}}\Omega\left(p^{l}\left|x_{p}-p^{l}a\right|_{p}\right)&\text{if }bp^{-r}-p^{l}a\in p^{-r}\mathbb{Z}_{p}\text{, }r\geq-l+1\\ &\\ 0&\text{if }bp^{-r}-p^{l}a\notin p^{-r}\mathbb{Z}_{p}\text{, }r\geq-l+1,\end{array}\right$.
where $\Omega\left(p^{l}\left|x_{p}-p^{l}a\right|_{p}\right)$ denotes the characteristic function of $p^{l}a+p^{l}\mathbb{Z}_{p}$. The above calculation has appeared in several publications; see, e.g., [23, Table 4.4].
Now, using (6.3), and (6.12),
(6.13) $\Omega\left(p^{l}\left|x_{p}-p^{l}a\right|_{p}\right)\Psi_{p}\left(x_{p},t\right)\Psi_{\infty}\left(\mathcal{M}\left(x_{p}\right),t\right)=$
${\displaystyle\sum\limits_{rbk}}{\displaystyle\sum\limits_{m}}a_{rbk,m}\left(t\right)\left\{\Omega\left(p^{l}\left|x_{p}-p^{l}a\right|_{p}\right)\right\}\psi_{rbk}\left(x_{p}\right)\theta_{m}\left(\mathcal{M}\left(x_{p}\right)\right)$
$={\displaystyle\sum\limits_{r<-l}}\ \ {\displaystyle\sum\limits_{b\in p^{r+l}b+p^{l+r}\mathbb{Z}_{p}}}\ {\displaystyle\sum\limits_{m}}a_{rbk,m}\left(t\right)\psi_{rbk}\left(x_{p}\right)\theta_{m}\left(\mathcal{M}\left(x_{p}\right)\right)+$
$\Omega\left(p^{l}\left|x_{p}-p^{l}a\right|_{p}\right)\left\{{\displaystyle\sum\limits_{r\geq-l+1}}\ {\displaystyle\sum\limits_{b\in p^{r+l}b+p^{l+r}\mathbb{Z}_{p}}}\ {\displaystyle\sum\limits_{m}}p^{\frac{-r}{2}}a_{rbk,m}\left(t\right)\theta_{m}\left(\mathcal{M}\left(x_{p}\right)\right)\right\}.$
6.2. The collapse mechanism
By a localized wavefunction, we mean  a compactly supported function in $L^{2}$. The existence of compactly supported wavefunctions $\Psi_{\infty}\left(\cdot,t\right)\in L^{2}\left(\mathbb{R}^{N}\right)$, $t\geq 0$, has a profound physical implications. For instance, the Dirac Hamiltonian does not admit localized wavefunctions: any wavefunction with positive energy has to be spread over all space ($\mathbb{R}^{3}$) at all times; see [56, Corollary 1.7]. The falsity of this result implies the violation of Einstein’s causality; see [56, Section 1.8.2].
We now discuss the localization (or collapse) of wavefunctions using the formulas (6.5) and (6.13). At the time of the measurement, the apparatus scans a small region of the microscopic space ($\mathbb{Q}_{p}$), which we assume to be a small ball $p^{l}a+p^{l}\mathbb{Z}_{p}$, where $l$ is a positive integer and $a\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. The ball $p^{l}a+p^{l}\mathbb{Z}_{p}$ corresponds to a small interval $\mathcal{M}\left(p^{l}a+p^{l}\mathbb{Z}_{p}\right)=\mathcal{M}(p^{l}a)+\left[0,p^{-l}\right]$, in the macroscopic space ($\mathbb{R}$).
The scanning causes a classical interaction between the apparatus $\mathcal{A}$ and the quantum object $\mathcal{S}$, which induces a localization of the wavefunction of $\mathcal{S}$, i.e., it passes from $\Psi_{p}\left(x_{p},t\right)$ to $\Psi_{p}\left(x_{p},t\right)\Omega\left(p^{l}\left|x-p^{l}a\right|_{2}\right)$, see (6.11). In turn, this localization induces a localization of the wavefunction of the system $\mathcal{S}+\mathcal{A}$ around a small interval, see (6.13). We can symbolically describe the localization of the wavefunction of the system as
$\Psi_{p}\left(x_{p},t\right)\text{ }\Psi_{\infty}\left(x_{\infty},t\right)\rightarrow\text{ \ }\\ \Omega\left(p^{l}\left|x-p^{l}a\right|_{2}\right)\Psi_{p}\left(x_{p},t\right)\text{ }\Psi_{\infty}\left(\mathcal{M}\left(x_{p}\right),t\right)=\Omega\left(p^{l}\left|x_{p}-p^{l}a\right|_{p}\right){\displaystyle\sum\limits_{m}}A_{m}\left(x_{p},t\right)\theta_{m}\left(\mathcal{M}\left(x_{p}\right)\right)\text{, }$
where
$A_{m}\left(x_{p},t\right)={\displaystyle\sum\limits_{r<-l}}\ \ {\displaystyle\sum\limits_{b\in p^{r+l}b+p^{l+r}\mathbb{Z}_{p}}}\ a_{rbk,m}\left(t\right)\psi_{rbk}\left(x_{p}\right)+{\displaystyle\sum\limits_{r\geq-l+1}}\ {\displaystyle\sum\limits_{b\in p^{r+l}b+p^{l+r}\mathbb{Z}_{p}}}\ p^{\frac{-r}{2}}a_{rbk,m}\left(t\right),$
for $x_{p}\in p^{l}a+p^{l}\mathbb{Z}_{p}$, see (6.13).
There are infinitely many Hamiltonians $\boldsymbol{H}_{p}$, which are (non-local) pseudo-differential operators, admitting eigenfunctions $\psi_{rbk}\left(x\right)$ supported in the ball $ap^{r}+p^{r}\mathbb{Z}_{p}$, for $a\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$, $r\in\mathbb{Z}$. Then, the proposed localization mechanism for the wavefunctions $\Psi_{p}\left(x_{p},t\right)$ $\Psi_{\infty}\left(x_{\infty},t\right)$ is mathematically realizable.
6.3. An additional remark
The collapse mechanism proposed does not work if we replace $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{Q}_{p}\right)$ by $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{R}\right)$. The calculations show in Table 6.12 use the fact that in $p$-adic topology, two balls are disjoint or one is contained in the other. Using that $\mathrm{supp}\psi_{rbk}=p^{-r}b+p^{-r}\mathbb{Z}_{p}$, the first line in Table 6.12 corresponds to the case $\mathrm{supp}\psi_{rbk}\subset B$; the second line corresponds to $B\subset\mathrm{supp}$ $\psi_{rbk}$, and the last line corresponds to $\mathrm{supp}$ $\psi_{rbk}\cap B=\varnothing$. This result is not valid in the standard topology of $\mathbb{R}$. Based on this fact, we assert that the proposed collapse mechanism is not valid in $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{R}\right)$.
6.4. A comparison with the GWR theory
The GRW theory posits that the collapse of the wavefunction occurs in position space, meaning that the wave function localizes to a specific region in space. The wavefunctions are physical entities that are subject to a spontaneous and random collapsing process in nature. The generalization of the GWR theory is the QM with spontaneous localization; this theory is an alternative to the standard QM.
In our proposal, the probability of interaction between the apparatus and the quantum particle (computed via the Born rule in the microscopic space-time $\mathbb{R}\times\mathbb{Q}_{p}$) has a density function compactly supported. We argue that this fact can be interpreted as that the wavefunction of the apparatus localizes during the measurement. From a mathematical perspective, the described localization process is similar to the one given in the GWR theory. However, in our framework, the wavefunctions are not physical entities; only the probability measures that they define have physical meaning. It is not necessary to introduce new physical constants, nor replace the Schrödinger equation with another equation involving nonlinear and stochastic terms. In our framework, the collapse of the wavefunction is a consequence of the difference between the geometry of the macroscopic realm and the geometry of the microscopic one.
It is believed that GWR theory is compatible with the experimental evidence of non-locality in the universe (the violation of Bell’s inequalities). Since the theory is formulated using the space-time model $\mathbb{R}\times\mathbb{R}^{3}$, the compatibility with special relativity is a significant problem. Finally, the quantum paradigm that the universe is not locally real demands that we choose between realism and non-locality. The bottom line question is whether GWR theory gives us a choice between (1) the universe is not local, or (2) the universe is real. We argue that the GWR theory cannot answer this question because it does not aim to explain the origin of non-locality in QM, nor does it model non-local processes in QM; instead, GWR theory uses non-locality only to explain the collapse of the wavefunction.
GRW theory establishes a natural collapse mechanism for wavefunctions, which does not require an external observer, as pointed out in the Copenhagen interpretation. Similarly, in our framework, the collapse of the wavefunction is a consequence of the geometry of the configuration space.
7. Two-slit Experiment in Quantum Mechanics
In this section, we apply the quantum formalism developed to the two-slit experiment. In [59], the authors provide a model of the two-slit experiment without using the wave properties of the particle: “Instead of a quantum wave passing through both slits, we have a localized particle with nonlocal interactions with the other slit.” In [22], the author proposes a $p$-adic model for the same experiment, giving a similar conclusion, where the non-local interactions are a consequence of the discreteness of the space $\mathbb{Q}_{p}$.
In [22], the evolution operators have the form $e^{-it\boldsymbol{D}^{\alpha}}$, $\alpha>0$, for $t\in\mathbb{R}$, where $\boldsymbol{D}^{\alpha}$ is non-local Taibleson-Vladimirov derivative; see (2.3). These operators are obtained from a Feller semigroup $e^{-t\boldsymbol{D}^{\alpha}}$, $t\in\mathbb{R}_{+}$, by a Wick rotation. The Feller feature means that there is a Markov process attached to the semigroup with state space $\mathbb{Q}_{p}$. Intuitively, this stochastic process describes a particle performing a random motion consisting of jumps between the points of a self-similar set; see [40, Chapter 2], [51, Chapter 2]. The free Schrödinger equation in natural units has the form $i\frac{\partial\Psi_{p}\left(x_{p},t\right)}{\partial t}=\boldsymbol{D}^{\alpha}\Psi\left(x_{p},t\right)$, $x_{p}\in\mathbb{Q}_{p},t\geq 0$. The function $\left|\Psi\left(x,t\right)\right|^{2}$ is a time-dependent probability density, and $\int_{B}\left|\Psi\left(x,t\right)\right|^{2}d^{N}x$ is the probability of finding a particle in the set $B$ at the time $t$.
The mathematical setup for the ($p$-adic) double-slit experiment is as follows. At time zero, there are two localized particles, which correspond to two localized Gaussian  probability densities at two different points $a$, $b\in\mathbb{Q}_{p}$, such $\left|a-b\right|_{p}>p^{-L}$. Geometrically, each slit corresponds to a ball centered at $a$ or $b$ with radius $p^{-L}$, $L\geq 0$, which is the size of each slit. These localized particles evolve in time under a non-local evolution operator $e^{-it\boldsymbol{D}^{\alpha}}$, $t\geq 0$; the non-locality is a consequence of the discreteness of space $\mathbb{Q}_{p}$. In standard QM, the fact that the de Broglie waves (the quantum waves) are considered physical waves explains the double-slit experiment as an interference phenomenon.
In the $p$-adic framework, de Broglie waves (‘the quantum waves’) are not physical entities; thus, the explanation of the double-slit experiment as an interference phenomenon is ruled out. However, the probability density $\left|\Psi_{p}\left(x_{p},t\right)\right|^{2}$ exhibits the classical interference patterns attributed to ‘quantum waves.’ In [22], we argue that the standard notion of trajectory, jointly with the fact that the de Broglie waves are just mathematical objects, concludes that in the double-slit experiment, each particle goes only through one of the slits.
The model for the two-slit experiment given here is based on the models given in [22] (case $\mathbb{Q}_{p}$), and [37] (case $\mathbb{R}$). We fix a non-zero rational number $s$, notice that $s\in\mathbb{Q}_{p}$ and $s\in\mathbb{R}$, and set
$\Psi_{p}^{\left(0\right)}\left(x_{p}\right)=A_{p}\left\{\Omega\left(p^{L}\left|x_{p}-s\right|_{p}\right)+\Omega\left(p^{L}\left|x_{p}+s\right|_{p}\right)\right\},$
where $L$ is a positive integer, and $A_{p}$ is a normalization constant such that $\left\|\Psi_{p}^{\left(0\right)}\left(x_{p}\right)\right\|_{2}=1$, and
$\Psi_{\infty}^{\left(0\right)}\left(x_{\infty}\right)=A_{\infty}\left\{e^{-\frac{\left(x-s\right)^{2}}{2\sigma^{2}}}+e^{-\frac{\left(x+s\right)^{2}}{2\sigma^{2}}}\right\},$
where $\sigma$ is a positive constant, and $A_{\infty}$ is a normalization constant such that $\left\|\Psi_{\infty}^{\left(0\right)}\left(x_{\infty}\right)\right\|_{2}=1$. Notice that $\Psi_{\infty}^{\left(0\right)}\left(x_{\infty}\right)$, $\Psi_{p}^{\left(0\right)}\left(x_{p}\right)$ are the superposition of two Gaussian packets located at $\pm s$. We propose the following model for the two-slit experiment:
(7.1) $\left\{\begin{array}[c]{l}i\frac{\partial}{\partial t}\Psi\left(x_{\infty},x_{p},t\right)=\left\{\frac{1}{2m_{p}}\boldsymbol{D}^{\alpha}-\frac{1}{2m_{\infty}}\frac{\partial^{2}}{\partial x_{\infty}}\right\}\Psi\left(x_{\infty},x_{p},t\right);\\ \\ \Psi\left(x_{\infty},x_{p},0\right)=\Psi_{p}^{\left(0\right)}\left(x_{p}\right)\Psi_{\infty}^{\left(0\right)}\left(x_{\infty}\right);\\ \\ \left\|\Psi\left(x_{\infty},x_{p},t\right)\right\|_{L^{2}\left(\mathbb{R}\times\mathbb{Q}_{p}\right)}=1\text{, }\end{array}\right$.
for $_{p}\in\mathbb{Q}_{p}, xx_{\infty}\in\mathbb{R}$, $t\geq 0$. This problem has a solution of the form $\Psi\left(x_{\infty},x_{p},t\right)=\Psi_{\infty}\left(x_{\infty},t\right)\Psi_{p}\left(x_{p},t\right)$, where
(7.2) $\left\{\begin{array}[c]{l}i\frac{\partial}{\partial t}\Psi_{\infty}\left(x_{\infty},t\right)=-\frac{1}{2m_{\infty}}\frac{\partial^{2}}{\partial x_{\infty}}\Psi_{\infty}\left(x_{\infty},t\right);\\ \\ \Psi_{\infty}\left(x_{\infty},0\right)=\Psi_{\infty}^{\left(0\right)}\left(x_{\infty}\right);\\ \\ \left\|\Psi\left(x_{\infty},t\right)\right\|_{L^{2}\left(\mathbb{R}\right)}=1\text{, }\end{array}\right$.
for $x_{\infty}\in\mathbb{R}$, $t\geq 0$, and
(7.3) $\left\{\begin{array}[c]{l}i\frac{\partial}{\partial t}\Psi_{p}\left(x_{p},t\right)=\frac{1}{2m_{p}}\boldsymbol{D}^{\alpha}\Psi_{p}\left(x_{p},t\right);\\ \\ \Psi_{p}\left(x_{p},0\right)=\Psi_{p}^{\left(0\right)}\left(x_{p}\right);\\ \\ \left\|\Psi_{p}\left(x_{p},t\right)\right\|_{L^{2}\left(\mathbb{Q}_{p}\right)}=1\text{, }\end{array}\right$.
where $x_{p}\in\mathbb{Q}_{p}$, $t\geq 0$. The model (7.2) provides a good approximation of the interference patterns for the two-slit experiment; see [37], and the references therein. The model (7.3) is a $p$-adic analogue of (7.2), which gives interference patterns similar to the standard ones, and provides a picture similar to the one given in [59]. Based on these reasons, we propose the model (7.1).
There are two different quantum states $\Psi_{\infty}\left(x_{\infty},t\right)$, $\Psi_{p}\left(x_{p},t\right)$ that produce two different interference patterns, one in the macroscopic realm $\left|\Psi_{\infty}\left(x_{\infty},t\right)\right|^{2}$, and the other $\left|\Psi_{p}\left(x_{p},t\right)\right|^{2}$ in the microscopic one. Here, it is important to remember that these two realms are connected via the map $\mathcal{M}:$ $\mathbb{Q}_{p}\rightarrow\mathbb{R}$. We argue that the pattern $\left|\Psi_{\infty}\left(x_{\infty},t\right)\right|^{2}$ can be detected by an apparatus located in $\mathbb$$\times R$}, while the pattern $\left|\Psi_{p}\left(x_{p},t\right)\right|^{2}$ remains hidden for the apparatus.
In [38], the authors propose that the classical interference patterns observed in experiments like the double-slit experiment originate from collective “bright” and “dark” states of light at the quantum level. The bright states are quantum states of light that can readily interact with a detector, while dark states are quantum states of light that cannot interact with the detector. In this framework, it is proposed that photons exist in these “dark” regions of the interference patterns. This proposal matches our previous discussion, if we interpret $\Psi_{\infty}\left(x_{\infty},t\right)$ as a bright state, and $\Psi_{p}\left(x_{p},t\right)$ as a dark state. However, from the perspective of the quantum formalism introduced here, we expect that the existence of dark states is a general fact in QM.
8. Conclusions and final discussion
QM on $\mathbb{C}^{N}$ is a particular case of $p$-adic QM, more precisely, it is QM on $L^{2}(\chi_{N}\left(\mathbb{Z}_{p}\right))$. $p$-Adic QM is non-local because the Hamiltonians (non-local operators) allow spooky action at a distance. Thus, this theory admits realism. The implementation of the space discreteness hypothesis requires using a completely disconnected space $\mathcal{X}$ as a model of the physical space at the microscopic level. In such a space, there are no continuous curves (world lines), so any theory using this type of space is not compatible with the theory of relativity. It does not imply any immediate prediction on the violation of Einstein’s causality.
We pick $\mathcal{X}=\mathbb{Q}_{p}$ because $p$-adic QM is the framework to formulate the non-locality tests and to formulate continuous-time quantum walks, which are a fundamental tool in quantum computing. This choice does not discard that other totally disconnected spaces can be more convenient in particular applications. Here, we use the $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{Q}_{p}\right)^{3}$ as a space-time model. Notice that passing from $\mathbb{R}\times\mathbb{R}^{3}$ to $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{Q}_{p}\right)^{3}$ demands passing from $4$-dimensions to $7$-dimensions. This larger space contains regions of the form $\mathbb{R}\times\mathbb{R}^{3}$, where relativistic QM is valid; this theory is local (uses local Hamiltonians) and real. At the macroscopic level, the realism is not in dispute. Also, the mentioned space contains regions of the form $\mathbb{R}\times\mathbb{Q}_{p}^{3}$, where $p$-adic QM is valid.
QM on the Hilbert space $L^{2}(\left(\mathbb{R}\times\mathbb{Q}_{p}\right)^{3})$, with evolution operators of the form $e^{it\left(\boldsymbol{H}_{p}+\boldsymbol{H}_{\infty}\right)}$ unifies the QM on $L^{2}(\mathbb{R}^{3})$ and $L^{2}(\mathbb{Q}_{p}^{3})$. In this framework, we propose a collapse mechanism for the wavefunction that gives an alternative solution to the measurement problem. In our proposal, the Schrödinger equation describes the evolution of the wavefunctions over time and their collapse.
The space-time model $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{Q}_{p}\right)^{3}$ allows breaking the causality and the Lorentz symmetry naturally, when passing from the macroscopic space $\mathbb{R}^{3}$ to the microscopic one $\mathbb{Q}_{p}^{3}$. The Lorentz symmetry is one of the most essential symmetries of the quantum field theory. While the validity of this symmetry continues to be verified with a high degree of precision [60], in the last thirty-five years, the experimental and theoretical studies of the Lorentz breaking symmetry have been an area of intense research, see, e.g., [61]- [62] and the references therein.
In [23], we introduced a $p$-adic Dirac equation that shares many properties with the standard one. In particular, the new equation also predicts the existence of pairs of particles and antiparticles and a charge conjugation symmetry. The geometry of space $\mathbb{Q}_{p}^{3}$ imposes substantial restrictions on the solutions of the $p$-adic Dirac equation,  in particular, it admits space-localized planes waves varying in time. This phenomenon does not occur in the standard case; see, e.g., [56, Section 1.8, Corollary 1.7]. On the other hand, we compute the transition probability from a localized state at time $t=0$ to another localized state at $t>0$, assuming that the space supports of the states are arbitrarily far away. It turns out that this transition probability is greater than zero for any time $t\in\left(0,\epsilon\right)$, for arbitrarily small $\epsilon$; see [23, Theorem 9.1]. Since this probability is nonzero for some arbitrarily small $t$, the system has a nonzero probability of getting between the mentioned localized states arbitrarily shortly, thereby propagating with superluminal speed in $\mathbb{R}\times\mathbb{Q}_{p}^{3}$.
Localization plays a central role in quantum field theory (QFT), and it is itself an area of intense research; see [63], and the references therein. Here, we point out that the geometry of space-time strongly influences localization. Using $\mathbb{R}\times\mathbb{R}^{3}$ as a model of space-time, Newton and Wigner studied the localization of particles in relativistic QM, [64]. The Newton-Wigner scheme in QFT predicts a phenomenon of superluminal spreading, [65], that contradicts the relativistic notion of causality. This result is a consequence of the Hegerfeldt theorem [66], [67], whose only hypotheses are the positivity of the energy of relativistic particles and the orthogonality condition of states localized in disjoint regions. The Newton-Wigner scheme contradicts the no-communication theorem; for this reason, it is not regarded as fundamental.
In 1988, Eberhard and Ross, using $\mathbb{R}\times\mathbb{R}^{3}$ as a space-time model, showed that the relativistic quantum field theory inherently forbids faster-than-light communication, [68]. This result is known as the no-communication theorem. It preserves the principle of causality in quantum mechanics and ensures that information transfer does not violate special relativity by exceeding the speed of light. It is relevant to mention that the no-communication theorem does not rule out the possible superluminal speed in $\mathbb{R}\times\mathbb{Q}_{p}^{3}$. Indeed, the study of the possible superluminal speed under the hypothesis of the space-time model $\mathbb{R}\times\left(\mathbb{R}\times\mathbb{Q}_{p}\right)^{3}$ is an open problem.
As we already pointed out in [21], the assumption that space discreteness implies the violation of Einstein’s causality does not immediately imply the possibility that one observer can transmit information to another observer exceeding the speed of light. The construction of a device that exploits the violation of the Einstein causality is an entirely different problem from the existence of the violation phenomena.

### v2

@@
 Abstract.
-[…]
+The space discreteness hypothesis asserts that the nature of space at short distances is radically different from that at large distances.
+Based on the Bronstein inequality, here, we use a totally disconnected topological space $\mathcal{X}$ as a model for the physical space.
+However, we consider the time as a real variable.
+In this framework, the Dirac-von Neumann formalism can be used.
+This discreteness hypothesis implies that given two different points in space, there is no continuous curve (a world line) joining them.
+Consequently, this hypothesis is not compatible with the theory of relativity.
+We propose $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{3}$ as a model of a space-time.
+For simplicity, we work out our models with $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)$ as the configuration space.
+Quantum mechanics (QM), in the sense of Dirac-von Neumann, on the Hilbert space $L^{2}\left(\mathbb{R}\times\mathcal{X}\right)$ is a non-local theory: the Hamiltonians are non-local operators, and thus, spooky action at a distance is allowed.
+The paradigm asserting that the universe is non-locally real implies that the proposed version of QM admits realism.
+This version of QM can be specialized to standard QM by using Hamiltonians acting on wavefunctions supported on the region $\mathbb$$\times R$}.
+We apply the developed formalism to the measurement problem.
+We propose a new mechanism for wavefunction collapse.
+The mechanism resembles that proposed by Ghirardi, Rimini, and Weber, but there are significant differences.
+The most important feature is that the Schrödinger equation describes the dynamics at all times, even at the moment of measurement.
+We also discuss a model for the two-slit experiment in which bright and dark states of light (recently proposed) naturally occur.
 1. Introduction
@@
 In the 1930s, Bronstein showed that general relativity and quantum mechanics imply that the uncertainty $\Delta x$ of any length measurement is greater than or equal to the Planck length, [7].
-The interpretation of this inequality is that there are no intervals, just points below the Planck length.
-We argue that this last assertion can be reformulated in precise mathematical form as: at the Planck length, the physical space can be modeled as a completely disconnected space $\mathcal{X}$.
+A widely accepted interpretation of this inequality is that there are no intervals, just points below the Planck length.
+We argue that this last assertion can be reformulated in precise mathematical terms as follows: at the Planck length, the physical space can be modeled as a totally disconnected space $\mathcal{X}$.
 In this framework, there is no continuous world line joining two different points in space $\mathcal{X}$; consequently, the assumption that $\mathcal{X}$ is a model of space is entirely incompatible with the theory of relativity.
@@
 A rigorous mathematical study of QM on Hilbert spaces of type $L^{2}\left(\left(\mathbb{R}\times\mathcal{X}\right)^{N}\right)$ is possible assuming suitable technical conditions, but this choice is quite involved.
-Here, we take $N=1$, $\mathbb{Q}_{p}$, the field of $p$-adic numbers; $\mathbb{Q}_{p}$ is a paramount example of a completely disconnected space, with a very rich mathematical structure.
+Here, we take $N=1$, $\mathcal{X}=\mathbb{Q}_{p}$, the field of $p$-adic numbers; $\mathbb{Q}_{p}$ is a paramount example of a totally disconnected space, with a very rich mathematical structure.
 In the following paragraphs, we provide an argument showing the physical relevance of the chosen model.
@@
 We divide the space-time into two realms: the macroscopic and the microscopic.
-In the first one, the model of the space-time is $\mathbb{R}\times\mathbb{R}^{3}$, while in the second one, it is $\mathbb{R}\times\mathcal{X}^{3}$, where $\mathcal{X}$ is a completely disconnected topological space.
+In the first one, the model of the space-time is $\mathbb{R}\times\mathbb{R}^{3}$, while in the second one, it is $\mathbb{R}\times\mathcal{X}^{3}$, where $\mathcal{X}$ is a totally disconnected topological space.
 The connectedness of a topological space is a topological property, which means it remains invariant under continuous transformations (homeomorphisms).
@@
 In this section, we fix the notation and collect some basic results on $p$-adic analysis that we use in this paper.
-For a detailed exposition on $p$-adic analysis, the reader may consult [15], [40], [41], [42].
-For a discussion on applications of $p$-adic analysis to physics, biology, computing, and other areas, the reader may consult [19], [22], [25], [43], [44], [45], [46], [48], and the references therein.
+For a detailed exposition on $p$-adic analysis, the reader may consult [15], [40]-[42].
+For a discussion on applications of $p$-adic analysis to physics, biology, computing, and other areas, the reader may consult [19], [22], [25], [43]-[46], [48], and the references therein.
 From now on, we use $p$ to denote a fixed prime number. Any non-zero $p$-adic number $x$ has a unique expansion of the form
@@
 As a topological space $\mathbb{Q}_{p}$ is homeomorphic to a Cantor-like subset of the real line, see, e.g., [15], [42].
-$\mathbb{Q}_{p}$ is a paramount example of a completely disconnected space, with a very rich mathematical structure.
+$\mathbb{Q}_{p}$ is a paramount example of a totally disconnected space, with a very rich mathematical structure.
 A function $\varphi:\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is called locally constant, if for any $a\in\mathbb{Q}_{p}$, there is an integer $l=l(a)$, such that
@@
 Here, it is relevant to recall that the locality of the Lagrangian densities is a required hypothesis in the formulation of functional integrals in field theory, [49, Section 5.1].
-Along this paper, we use discrete space as a synonymous of completely disconnected space.
+Along this paper, we use discrete space as a synonymous of totally disconnected space.
 The non-locality is natural in discrete spaces like $\mathbb{Q}_{p}$.
-On a completely disconnected topological space, like $\mathbb{Q}_{p}$, the locally constant functions play a central role.
+On a totally disconnected topological space, like $\mathbb{Q}_{p}$, the locally constant functions play a central role.
 A function $\varphi:\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is locally constant, if $\varphi\left(x+x^{\prime}\right)-\varphi\left(x\right)=0$ for $x^{\prime}$ sufficiently small.
@@
 We denote by $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ the $N$-dimensional Hilbert space spanned by (3.1).
-Then, $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ is isomorphic to $\mathbb{C}^{N}$, as Hilbert spaces.
+Then, $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ and $\mathbb{C}^{N}$ are isomorphic as Hilbert spaces.
 Notice that the isomorphism exists for any $p$ and $l$ such that $N\leq p^{l}$.
-The advantage of using $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ over $\mathbb{C}^{N}$ is that we can construct, in a natural way, continuos versions of quantum models in $\mathbb{C}^{N}$.
+The advantage of using $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ over $\mathbb{C}^{N}$ is that we can construct, in a natural way, continuous versions of quantum models in $\mathbb{C}^{N}$.
 We identify $p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ with $e_{I}$, where $\left\{e_{I}\right\}_{I\in G_{l}^{0}}$ is the canonical basis of $\mathbb{C}^{N}$.
@@
 $i\frac{\partial}{\partial t}\Psi\left(x,t\right)=\boldsymbol{H}\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p}\text{, }t\geq 0,$
-is a continuous Schrödinger equation attached to matrix $\left[H_{J,K}\right]_{1\leq J,K\leq N}$. For further details, the reader may consult [21]; see also [24], [25], and the references therein.
+is a continuous Schrödinger equation attached to matrix $\left[H_{J,K}\right]_{1\leq J,K\leq N}$. For further details, the reader may consult [21]; see also [24]-[25], and the references therein.
 In 1964, Bell established that quantum mechanics is inherently non-local through a non-locality test known now as Bell inequalities, [2].
@@
 4.2. Some results about tensor product of Hilbert spaces
-We set $L^{2}(\mathbb{R},dx_{\infty})=L^{2}(\mathbb{R})$, where  denotes the Lebesgue measure, and $L^{2}\left(\mathbb{Q}_{p},dx_{p}\right)=L^{2}\left(\mathbb{Q}_{p}\right)$, where $dx_{p}$ denotes the Haar measure of $\mathbb{Q}_{p}$.
+We set $L^{2}(\mathbb{R},dx_{\infty})=L^{2}(\mathbb{R})$, where $dx_{\infty}$ denotes the Lebesgue measure, and $L^{2}\left(\mathbb{Q}_{p},dx_{p}\right)=L^{2}\left(\mathbb{Q}_{p}\right)$, where $dx_{p}$ denotes the Haar measure of $\mathbb{Q}_{p}$.
 We denote by $dx_{\infty}dx_{p}$, the product measure of $dx_{\infty}$ and $dx_{p}$ on $\mathbb{R}\times\mathbb{Q}_{p}$.
@@
 for some fixed $\alpha\in\mathbb{Q}_{p}$, where the standard (local) QM is valid, and regions of the form
-$\mathbb{R}\times\left(\left\{\beta\right\}\times\mathbb{Q}_{p}\right)^{N}\simeq\mathbb{R}\times\mathbb{Q}_{p}^{3N},$
+$\mathbb{R}\times\left(\left\{\beta\right\}\times\mathbb{Q}_{p}\right)^{N}\simeq\mathbb{R}\times\mathbb{Q}_{p}^{N},$
 for some fixed $\beta\in\mathbb{R}$, where $p$-adic (non-local) QM is valid.
@@
 In this section, we review quickly the measurement problem and the Ghirardi-Rimini-Weber (GRW) collapse theory.
-In the next section, we propose a solution to the measurement problem; the collapse mechanism resembles the one in the GRW theory, but there are several fundamental differences.
+In the next section, we propose a solution to the measurement problem.
+The new collapse mechanism resembles the one in the GRW theory, but there are several fundamental differences.
 5.1. The measurement problem
@@
 $P(r)=\sqrt{{\displaystyle\int\limits_{\mathbb{R}}}\left|g_{r}\left(x\right)\Psi\left(x,t_{-}\right)\right|^{2}dx}.$
-The parameter $r$ is random with a probability distribution $P(r)$.
+The parameter $r$ is a random variable with a probability distribution $P(r)$.
 The localization (5.2) occurs at random times, with a Poisson distribution with mean frequency $\lambda$.
@@
 Thus, this theory admits realism.
-The implementation of the space discreteness hypothesis requires using a completely disconnected space $\mathcal{X}$ as a model of the physical space at the microscopic level.
+The implementation of the space discreteness hypothesis requires using a totally disconnected space $\mathcal{X}$ as a model of the physical space at the microscopic level.
 In such a space, there are no continuous curves (world lines), so any theory using this type of space is not compatible with the theory of relativity.
@@
 The construction of a device that exploits the violation of the Einstein causality is an entirely different problem from the existence of the violation phenomena.
+Finally, we mention that in adelic QM, the configuration space is the product $\mathbb{R}\times\mathbb{A}_{\mathbb{Q}}$, where $\mathbb{A}_{\mathbb{Q}}$ is the ring of finite adeles, which consists of sequences $\left\{x_{p}\right\}_{p\geq 2}$ from ${\textstyle\prod\nolimits_{p\geq 2}}\mathbb{Q}_{p}$ such that $x_{p}\in\mathbb{Z}_{p}$ for $p$ sufficiently large.
+Dragovich and his collaborators have studied several quantum models on the infinite-dimensional space $\mathbb{R}\times\mathbb{A}_{\mathbb{Q}}$; see, e.g., [69]-[71].
+Extra dimensions in physical models represent a nontrivial feature.
+Our solution to the measurement problem relies on extra dimensions.
+As we already noted, the $p$-adic QM describes continuous-time quantum walks, including those that stem from graphs.
+The physical implementation of these random walks generally involves particles or excitations on networks of size scales ranging from a few nanometers to micrometers.
+Therefore, it naturally leads the author to ask whether extra dimensions exist at the nanoscale.
+This fact matches Vafa’s proposal for the existence of at least one extra spatial dimension at the nanoscale, [72].
+Data availability statement
+No new data were created or analyzed in this study. Data sharing is not applicable to this paper.